# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ IV.1.1 Signal enhancement
We now present the numerical results for the readout mean rate enhancement $\Delta{r_{\rm R}}:={r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}$ versus ${\tau_{\rm c}}$, for two values of the fraction of shared noise ${\alpha_{\rm V}}$ (see Appendix C for the numerical range of probed parameters, and Appendix F for different choices of such parameters).
@@
First, we observe that introducing noise correlations enhances the firing rate difference for all the considered values of ${\tau_{\rm c}}$ (Fig. 8-A). This effect is not simply equivalent to the LIF neuron firing rate enhancement in the presence of noise correlations (see Fig. 4), but it correspond to an increase of the mean firing rate difference $\Delta{r_{\rm R}}$, in the presence of noise correlations.
@@
For low enough ${\tau_{\rm c}}$ (and constant ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$), the colored-noise amplitude ${\sigma_{\rm c}}$ vanishes, so that the numerical $\Delta{r_{\rm R}}$ is consistent with that of the white noise LIF $\Delta{r_{\rm w}}$. For larger but low enough values of ${\tau_{\rm c}}$, the Moreno perturbative approximation to ${r_{\rm R}}$ for low ${\tau_{\rm c}}/{\tau_{\rm m}}$ holds as expected. For large values of ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, we observe a disagreement between the numerical estimation and the Moreno approximation, and an agreement between the numerical estimation and the quenched-noise approximation, as expected (see Sec. III.2 and Appendix F)  [^2] Notice that both the Moreno solution and the quenched-noise solution exhibit a constant behaviour in ${\tau_{\rm c}}$ for large ${\tau_{\rm c}}$ (and constant ${v_{\rm c}}$), as expected from Eqs. (9,10).. For intermediate values of ${\tau_{\rm c}}\lesssim{\tau_{\rm m}}$, although outside the validity limits of both approximations, we observe a nice agreement between the numerical $\Delta{r_{\rm R}}$ and the interpolation of both Moreno solutions [17] (although this agreement is sensitive to the interpolating time, here set to $t_{\rm in}=2.5{\tau_{\rm m}}$).
@@
Thus, for fixed values of the parameters ${\mu_{\rm I}}^{(\pm)},{\tau_{\rm m}},w$, increasing noise correlations leads to an increase in the readout neuron firing rate difference $\Delta{r_{\rm R}}$, quantitatively understood within the Moreno et al solution for low ${\tau_{\rm c}}$ and within the quenched-noise approximation (Sec. III.2) for large enough ${\tau_{\rm c}}$.
@@
Figure 8: Mean rate enhancement, rate standard deviation and SNR, as a function of ${\tau_{\rm c}}$. (A) Readout mean rate increment $\Delta{r_{\rm R}}={r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}$ vs ${\tau_{\rm c}}$. Different curves/colors in each panel correspond to both values of ${\alpha_{\rm V}}=0,0.9$. Void points correspond to the infinite-NIS extrapolated procedure (see Appendix C) (and the error-bars indicate the uncertainty of the regression), while the semi-transparent points correspond to the raw numerical estimations $\tilde{r_{\rm R}}$ for the highest NIS (and the error-bars are $(\tilde{v_{\rm R}}{T_{\rm c}})^{1/2}$). The dashed lines are the interpolated Moreno perturbative solution in [17], using an interpolating time $t_{\rm in}=2.5{\tau_{\rm m}}$. The continuous line (shown for ${\tau_{\rm c}}>{\tau_{\rm m}}$ only) is the quenched-noise value, Eq. (9). The vertical dotted line signals the value of ${\tau_{\rm m}}$. (B): Standard deviation of the stimulus-averaged readout firing rate in the unit time interval, $[({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2]^{1/2}{T_{\rm c}}^{1/2}$ vs ${\tau_{\rm c}}$. Here, only the raw estimations of the variance are shown (filled points). The error-bars are the $\chi^{2}$ confidence interval with a probability $\alpha_{\chi}=0.95$. The pairs of dashed lines (shown for ${\tau_{\rm c}}>{\tau_{\rm m}}$ only) correspond to the lower and upper bounds for the variance in the quenched-noise approximation, (13). We attribute the slight disagreement between the numerical estimations and ${v_{\rm q}}$ to a finite-NIS underestimation of the numerical variance. (C): Readout SNR in the unit time interval (of which panels A and B are, respectively, the numerator and the denonminator), vs ${\tau_{\rm c}}$. The numerical points correspond to the raw estimators for both the numerator and the denominator. The thin dotted lines are the encoding signal-to-noise ratio in the unit time interval, Eq. (8). The simulation parameters are: $(\mu_{I}^{(+)},\mu_{I}^{(-)})=(32.65,41.5)$, $w=0.45$, ${v_{\rm V}}=9.80$.
IV.1.2 Rate variance enhancement
@@
We now consider the denominator of the SNR, i.e. the stimulus-averaged variance of the readout firing rate, $({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2$. First, for low enough ${\tau_{\rm c}}$ we observe an agreement between $\tilde{v_{\rm R}}$ and the Poisson-Moreno approximation (Fig. 8-B). We attribute small deviations to a numerical underestimation of the variance due to the finite-NIS effect (Appendix K). We notice that in this case we do not employ the NIS-corrected estimation for the rate variance, since it exhibits larger statistical errors. Second, the increment of $\tilde{v_{\rm R}}$ for large values of ${\tau_{\rm c}}$ in the presence of noise correlations can be explained, within the quenched-noise approximation, as the result of the integration of the over-Poissonian part of the temporal correlation function ${g_{\rm q}}$ (see Eq. (13), and Appendix H for a further, intuitive explanation of the enhancement of $\tilde{v_{\rm R}}$ for large ${\tau_{\rm c}}$).
@@
Overall, the Poisson-Moreno and quenched-disorder approximations for ${v_{\rm R}}$ describe the numerical $\tilde{v_{\rm R}}$ in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ regimes, respectively. For large ${\tau_{\rm c}}$ the variance ${v_{\rm R}}$ increases with ${\tau_{\rm c}}$ (see the second term in Eq. (11) and Eq. (13)), while the mean rate ${r_{\rm R}}$ stays constant with ${\tau_{\rm c}}$ (see Eq. (10)).
IV.1.3 SNR enhancement
@@
From the results of the previous sections, it follows that the SNR exhibits a non-monotonic behaviour with ${\tau_{\rm c}}$ (Fig. 8-C). For large enough ${\tau_{\rm c}}$, noise correlations lower the readout SNR below its value in absence of noise correlations. For high values of ${\tau_{\rm c}}$, the increment of the SNR numerator, $\Delta{r_{\rm R}}$, depends only on ${v_{\rm c}}$, so that it stays constant in ${\tau_{\rm c}}$ (see Fig. 8-A and Sec. III.2). Instead, the rate variance increases with ${\tau_{\rm c}}$ due to the over-Poissonian correction (which increases with both ${\tau_{\rm c}}$ and ${v_{\rm c}}$, see Eq. (13a) and Appendix H). Beyond a certain value of ${\tau_{\rm c}}$, the increment in the denominator becomes prevalent, thus leading to a decrement of the SNR. For sufficiently small ${\tau_{\rm c}}$, when the enhancement in the rate variance $({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2$ is not enough to overcome that of $\Delta{r_{\rm R}}$ in the SNR numerator, we observe that the noise correlations induce an enhancement of the SNR. This constitutes the apparently paradoxical effect according to which encoding correlations that lower the encoding-SNR enhance the readout SNR.
@@
In Fig. 8-C we report as well the encoding SNR (see Eq. (8)), which, by construction, decreases with the noise correlation amplitude, and which may actually become lower than the readout SNR (please, see a discussion on this point in Appendix I).
@@
For details regarding the small theoretical-numerical disagreement for low ${\tau_{\rm c}}$ in Fig. 8-C, we refer to the Appendix K, where we attribute it to the finite-NIS effect.
@@
In sum, the SNR presents a non-monotonic behavior with ${\tau_{\rm c}}$, that is quantitatively explained by the quenched-noise approximation (and, for low ${\tau_{\rm c}}$, by Moreno et al approximation to $o({\tau_{\rm c}}/{\tau_{\rm m}})$).
IV.2 The conditions for enhanced-by-consistency SNR to occur
@@
In this section we address for which general values of the encoding and decoding parameters we expect the SNR enhancement to occur in the presence of noise correlations. The quenched-noise approximation for the rate mean and variance allows us to draw necessary conditions on the NERM parameters, at least for sufficiently large ${\tau_{\rm c}}$. Our conclusions are drawn from the behavior of ${r_{\rm q}},{v_{\rm q}}$ as a function of ${\mu_{\rm I}}$ and ${\sigma_{\rm c}}$ (Fig. 9). In this section, we use the quenched-noise approximation and, in particular, the hand-waving expression for the rate variance, Eq. (72) in Appendix G, which is already a good approximation for the behavior of the rate variance (Fig. 7).
@@
Conditions on ${\mu_{\rm I}}^{(\pm)}$. We said that, for the colored noise to enhance the firing rate ${r_{\rm R}}$, the value of ${\mu_{\rm I}}$ should be safely lower than the inflection point of the LIF (circle in Fig. 9-A, see Sec. III.3). Given the parameters of the white LIF ${\tau_{\rm m}},{\sigma_{\rm I}}$, the inflection point occurs at the value of ${\mu_{\rm I}}$ for which $g_{1}^{2}+g_{1}^{\prime}=0$, where $g_{1}={r_{\rm w}}^{\prime}/{r_{\rm w}}$. The inflection point ${\mu_{\rm I}}^{*}$ of ${r_{\rm w}}({\mu_{\rm I}})$ is also the value of ${\mu_{\rm I}}$ below which ${r_{\rm q}}>{r_{\rm w}}$ for low enough ${v_{\rm c}}$ (indeed, expanding ${r_{\rm q}}$ in (10) to $o({v_{\rm c}})$, we get that ${r_{\rm q}}\geq{r_{\rm w}}$ for ${\mu_{\rm I}}\leq{\mu_{\rm I}}^{*}$, v. Fig. 9). Since the SNR numerator is $\Delta{r_{\rm R}}\leq{r_{\rm R}}({\mu_{\rm I}}^{(+)})$, a necessary condition for the SNR to be positive is ${\mu_{\rm I}}^{(+)}<{\mu_{\rm I}}^{*}$. We can make this condition more stringent by imposing that, beyond the numerator, the SNR itself should increase. The numerator ${r_{\rm R}}^{(+)}$ and the denominator ${v_{\rm R}}^{(+)}$ both increase with ${\sigma_{\rm c}}$, so that both effects cancel at first order in ${v_{\rm c}}$. In fact, to order $o({v_{\rm c}})$, the condition $\text{SNR}_{\rm q}=\text{SNR}_{\rm w}={r_{\rm w}}^{1/2}$ translates into $g_{1}^{\prime}+g_{1}^{2}=2{\tau_{\rm c}}{r_{\rm w}}g_{1}^{2}$, implying that the value of ${\mu_{\rm I}}$, say $\hat{\mu_{\rm I}}$ (green line in Fig. 9), for which this relation is satisfied does not depend on ${v_{\rm c}}$. Indeed, the ${\mu_{\rm I}}$ value at which the SNR becomes positive depends slowly on ${\sigma_{\rm c}}$ (Fig. 9). Therefore, the condition, ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$ can be taken as a necessary condition on ${\mu_{\rm I}}^{(+)}$ for the SNR to increase in the presence of noise correlations, in the NERM (and for large enough ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$). This necessary condition is, of course, not sufficient to guarantee that $\text{SNR}_{\rm q}>\text{SNR}_{\rm w}$, since so far we have neglected ${r_{\rm R}}^{(-)}$ in front of ${r_{\rm R}}^{(+)}$. Non-negligible values of ${r_{\rm w}}({\mu_{\rm I}}^{(-)})$ (hence, of ${r_{\rm R}}^{(-)}$) may change this scenario (Fig. 21 in Appendix L). The larger ${\mu_{\rm I}}^{(-)}$, the lower $\Delta{r_{\rm R}}$ and, for low values of ${\mu_{\rm I}}^{(-)}$, the larger ${v_{\rm R}}^{(-)}$ as well.
@@
For larger values of ${\sigma_{\rm c}}$, the (leftmost) inflection points of ${r_{\rm w}}$ and ${v_{\rm R}}$ occur at lower values of ${\mu_{\rm I}}$ (Fig. 9), implying that ${\mu_{\rm I}}^{(-)}$ should take even lower values for the SNR to increase. For large enough values of ${\sigma_{\rm c}}$ (or ${v_{\rm c}}$), the SNR denominator increases faster with ${\sigma_{\rm c}}$ than the numerator, because of the over-Poissonian correction to the variance, cf. the second term in Eq. (13a), or Eq. (72) in Appendix H. This is the origin of the non-monotonous dependence of the SNR on the noise amplitude.
@@
In sum, we derived a set of necessary conditions on ${\mu_{\rm I}}^{(\pm)}$ for the SNR to increase, for large enough ${\tau_{\rm c}}$, and when the rest of the variables are kept constant, is: (c1) ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$, and (c2) that ${\mu_{\rm I}}^{(-)}<\hat{\mu_{\rm I}}^{(-)}$ is low enough, with the upper bound $\hat{\mu_{\rm I}}^{(-)}$ being a decreasing function of ${\sigma_{\rm c}}$. The larger the amplitude of noise correlations, the larger should be the length of the current gap $\Delta{\mu_{\rm I}}$ for the SNR to increase with noise-correlations.
@@
Figure 9: Readout rate mean, variance and excess of SNR for negligible values of ${r_{\rm R}}^{(-)}$. (A) Mean firing rate ${r_{\rm R}}({\mu_{\rm I}})$ of the white-noise LIF (dashed line), and the quenched-noise colored-noise LIF ${r_{\rm q}}({\mu_{\rm I}})$ (continuous lines) for various values of ${\sigma_{\rm c}}$ (or ${v_{\rm c}}={\sigma_{\rm c}}^{2}/({\tau_{\rm c}})$, different curves). For each ${\mu_{\rm I}}$, the value of ${\sigma_{\rm I}}$ is set to ${\sigma_{\rm I}}=(w{\mu_{\rm I}})^{1/2}$, with $w=10^{-3}$. As in Fig. 5, the circle and error-bars indicate the inflection point and ${v_{\rm c}}^{1/2}$, while the black dotted line indicate the activation function of the deterministic LIF $r_{0}$. (B) A zoom of the panel A, in which we report as well the quenched-noise rate standard deviation, ${v_{\rm q}}^{1/2}$ (dashed-dotted lines). The vertical lines indicate the maximum of each ${r_{\rm q}}({\mu_{\rm I}})$ curve. (C) The difference between the SNR $[{r_{\rm q}}({\mu_{\rm I}})-{r_{\rm q}}({\mu_{\rm I}}^{(-)})]/[({v_{\rm q}}({\mu_{\rm I}})+{v_{\rm q}}({\mu_{\rm I}}^{(-)})/2]^{1/2}$ (taking a low ${\mu_{\rm I}}^{(-)}=260$, actually equivalent to neglect ${r_{\rm q}}^{(-)}$) in the presence of colored noise, minus the white-noise SNR, $=[{r_{\rm w}}({\mu_{\rm I}})-{r_{\rm w}}({\mu_{\rm I}}^{(-)})]/[({v_{\rm w}}({\mu_{\rm I}})+{v_{\rm w}}({\mu_{\rm I}}^{(-)}))/2]^{1/2}$, assuming a Poissonian rate variance, ${v_{\rm w}}={r_{\rm w}}$. The vertical dashed line corresponds to our upper bound $\hat{\mu_{\rm I}}$. Rates and variances are calculated with a ${\sigma_{\rm I}}^{2}=w{\mu_{\rm I}}$ that varies with ${\mu_{\rm I}}$ for fixed $w$ (see the repository [48]). The rest of the model parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.0025$, $w=0.001$, ${\mu_{\rm I}}^{(-)}=260$.
Conditions on other variables. These considerations are enough to understand the qualitative dependence of the SNR with respect to the rest of the LIF parameters. We exemplify such dependencies in Fig. 10 and in Figs. 23,24,25 in Appendix M.

### v2

@@ IV.2.1 Signal enhancement
We now present the numerical results for the readout mean rate enhancement $\Delta{r_{\rm R}}:={r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}$ versus ${\tau_{\rm c}}$, for two values of the fraction of shared noise ${\alpha_{\rm V}}$. We refer to Appendix C for the numerical range of probed parameters, and to Appendix G for different choices of such parameters.
@@
First, we observe that introducing noise correlations enhances the firing rate difference for all the considered values of ${\tau_{\rm c}}$ (Fig. 9-A). This effect is not simply equivalent to the LIF neuron firing rate enhancement in the presence of noise correlations (see Fig. 5), but it corresponds to an increase of the mean firing rate difference $\Delta{r_{\rm R}}$, in the presence of noise correlations.
@@
For low enough ${\tau_{\rm c}}$ (and constant ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$), the colored-noise amplitude ${\sigma_{\rm c}}$ vanishes, so that the numerical $\Delta{r_{\rm R}}$ is consistent with that of the white noise LIF $\Delta{r_{\rm w}}$. For larger but low enough values of ${\tau_{\rm c}}$, the Moreno perturbative approximation to ${r_{\rm R}}$ for low ${\tau_{\rm c}}/{\tau_{\rm m}}$ holds as expected. For large values of ${\tau_{\rm c}}\gg{\tau_{\rm m}}$, we observe a disagreement between the numerical estimation and the Moreno approximation, and an agreement between the numerical estimation and the quenched-noise approximation, as expected (see Sec. III.2 and Appendix G)  [^2] Notice that both the Moreno solution and the quenched-noise solution exhibit a constant behavior in ${\tau_{\rm c}}$ for large ${\tau_{\rm c}}$ (and constant ${v_{\rm c}}$), as expected from Eqs. (9,10).. For intermediate values of ${\tau_{\rm c}}\lesssim{\tau_{\rm m}}$, although outside the validity limits of both approximations, we observe a nice agreement between the numerical $\Delta{r_{\rm R}}$ and the interpolation of both Moreno solutions [17] (although this agreement is sensitive to the interpolating time, here set to $t_{\rm in}=2.5{\tau_{\rm m}}$).
@@
Thus, for fixed values of the parameters ${\mu_{\rm I}}^{(\pm)},{\tau_{\rm m}},w$, increasing noise correlations leads to an increase in the readout neuron firing rate difference $\Delta{r_{\rm R}}$, quantitatively understood within the Moreno et al solution for low ${\tau_{\rm c}}$ and within the quenched-noise approximation for large enough ${\tau_{\rm c}}$.
@@
Figure 9: Mean rate enhancement, rate standard deviation and SNR, as a function of ${\tau_{\rm c}}$. (A) Readout mean rate increment $\Delta{r_{\rm R}}={r_{\rm R}}^{(+)}-{r_{\rm R}}^{(-)}$ vs ${\tau_{\rm c}}$. Different curves/colors in each panel correspond to both values of ${\alpha_{\rm V}}=0,0.9$. Stars correspond to the infinite-NIS extrapolated procedure (see Appendix C) (and the error-bars indicate the uncertainty of the regression), while the semi-transparent squares correspond to the raw numerical estimations $\tilde{r_{\rm R}}$ for the highest NIS (and the error-bars are $(\tilde{v_{\rm R}}{T_{\rm c}})^{1/2}$). The dotted lines are the interpolated Moreno perturbative solution in [17], using an interpolating time $t_{\rm in}=2.5{\tau_{\rm m}}$. The dashed line (shown for ${\tau_{\rm c}}>{\tau_{\rm m}}$ only) is the quenched-noise analytical expression, Eq. (10). The vertical dotted line signals the value of ${\tau_{\rm m}}$. (B): Standard deviation of the stimulus-averaged readout firing rate in the unit time interval, $[({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2]^{1/2}{T_{\rm c}}^{1/2}$ vs ${\tau_{\rm c}}$. Here, only the raw estimations of the variance are shown. The error-bars are the $\chi^{2}$ confidence interval with a probability $\alpha_{\chi}=0.95$. The pairs of dashed lines (shown for ${\tau_{\rm c}}>{\tau_{\rm m}}$ only) correspond to the lower and upper bounds for the variance in the quenched-noise approximation, Eq. (13). We attribute the slight disagreement between the numerical estimations and ${v_{\rm q}}$ to a finite-NIS underestimation of the numerical variance. (C): Readout SNR in the unit time interval (of which panels A and B are, respectively, the numerator and the denonminator), vs ${\tau_{\rm c}}$. The numerical points correspond to the raw estimators for both the numerator and the denominator. We remind that ${\sigma_{\rm c}}$ varies with ${\tau_{\rm c}}$ according to Eq. (6c). The simulation parameters are: $(\mu_{I}^{(+)},\mu_{I}^{(-)})=(32.65,41.5)$, $w=0.45$, ${N}=20$, ${v_{\rm V}}=9.8$, ${\tau_{\rm m}}=0.005$.
IV.2.2 Rate variance enhancement
@@
We now consider the denominator of the SNR, i.e. the stimulus-averaged variance of the readout firing rate, $({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2$. First, for low enough ${\tau_{\rm c}}$ we observe an agreement between $\tilde{v_{\rm R}}$ and the Poisson-Moreno approximation (Fig. 9-B). We attribute the small deviations to a numerical underestimation of the variance due to the finite-NIS effect (see Appendix M). We notice that in this case we do not employ the NIS-corrected estimation for the rate variance, since it exhibits larger statistical errors. Second, the over-Poissonian increment of $\tilde{v_{\rm R}}$ for large values of ${\tau_{\rm c}}$ in the presence of noise correlations can only be explained, within the quenched-noise approximation, as the result of the integration of the temporal correlation function ${g_{\rm q}}$ (see Eq. (13), and Appendix I for a further, intuitive explanation of the enhancement of $\tilde{v_{\rm R}}$ for large ${\tau_{\rm c}}$).
@@
Overall, both approximations, the Poisson-Moreno and the quenched-noise approximations describe the numerical behavior of $\tilde{v_{\rm R}}$ in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ regimes, respectively. For large ${\tau_{\rm c}}$, the variance ${v_{\rm R}}$ increases with ${\tau_{\rm c}}$ (see the second term in Eq. (11) and Eq. (13)), while the mean rate ${r_{\rm R}}$ stays constant with ${\tau_{\rm c}}$ (as expected, see Eq. (10)).
IV.2.3 SNR enhancement
@@
From the results of the previous subsections, it follows that the SNR exhibits a non-monotonic behavior with ${\tau_{\rm c}}$ (Fig. 9-C). For large enough ${\tau_{\rm c}}$, noise correlations lower the readout SNR below its value in absence of noise correlations. For high values of ${\tau_{\rm c}}$, the increment of the SNR numerator, $\Delta{r_{\rm R}}$, depends only on ${v_{\rm c}}$, so that it stays constant in ${\tau_{\rm c}}$ (see Fig. 9-A and Sec. III.2). Instead, the rate variance increases with ${\tau_{\rm c}}$ due to the over-Poissonian correction (which increases with both ${\tau_{\rm c}}$ and ${v_{\rm c}}$, see Eq. (13a) and Appendix I). Beyond a certain value of ${\tau_{\rm c}}$, the increment in the denominator becomes prevalent, thus leading to a decrement of the SNR. For sufficiently small ${\tau_{\rm c}}$, when the enhancement in the rate variance $({v_{\rm R}}^{(+)}+{v_{\rm R}}^{(-)})/2$ is not enough to overcome that of $\Delta{r_{\rm R}}$ in the SNR numerator, we observe that the noise correlations induce an enhancement of the SNR. This constitutes the apparently paradoxical effect according to which encoding correlations that lower the encoding-SNR, enhance the readout SNR.
@@
We have also compared the readout SNR with the encoding SNR (see Eq. (8)) (see Fig. 27 in Appendix G and Fig. 29 in Appendix M). By construction, the encoding SNR decreases with the noise correlation amplitude, and it may actually become lower than the readout SNR (please, see a discussion on this point in Appendix K).
@@
For details regarding the small theoretical-numerical disagreement for low ${\tau_{\rm c}}$ in Fig. 9-C, we refer to the Appendix M, where we attribute it to the finite-NIS effect.
@@
In sum, the SNR presents a non-monotonic behavior with ${\tau_{\rm c}}$, that is quantitatively explained by the quenched-noise approximation.
IV.3 The conditions for enhanced-by-consistency SNR to occur
@@
In this section we address for which general values of the encoding and decoding parameters we expect the SNR enhancement to occur in the presence of noise correlations. The quenched-noise approximation for the rate mean and variance allows us to draw necessary conditions on the NERM parameters, at least for sufficiently large ${\tau_{\rm c}}$. Our conclusions are drawn from the behavior of ${r_{\rm q}},{v_{\rm q}}$ as a function of ${\mu_{\rm I}}$ and ${\sigma_{\rm c}}$ (Fig. 10). In this section, we use the quenched-noise approximation and, in particular, the hand-waving expression for the rate variance, Eq. (110) in Appendix H, which is already a good approximation for the behavior of the rate variance (Fig. 8).
@@
Conditions on ${\mu_{\rm I}}^{(\pm)}$ and dependence on ${\sigma_{\rm c}}$. In Sec. III.3 we said that, for the colored noise to enhance the firing rate ${r_{\rm R}}$, the value of ${\mu_{\rm I}}$ should be safely lower than the inflection point of the LIF (circle in Fig. 10-A). Given the parameters of the white LIF ${\tau_{\rm m}},{\sigma_{\rm I}}$, the inflection point occurs at the value of ${\mu_{\rm I}}$ for which $g_{1}^{2}+g_{1}^{\prime}=0$, where $g_{1}={r_{\rm w}}^{\prime}/{r_{\rm w}}$. The inflection point ${\mu_{\rm I}}^{*}$ of ${r_{\rm w}}({\mu_{\rm I}})$ is also the value of ${\mu_{\rm I}}$ below which ${r_{\rm q}}>{r_{\rm w}}$ for low enough ${v_{\rm c}}$ (indeed, expanding ${r_{\rm q}}$ in (10) to $o({v_{\rm c}})$, we get that ${r_{\rm q}}\geq{r_{\rm w}}$ for ${\mu_{\rm I}}\leq{\mu_{\rm I}}^{*}$, v. Fig. 10). We can proof (see Appendix J) that the condition ${\mu_{\rm I}}^{(+)}<{\mu_{\rm I}}^{*}$ is actually a necessary condition for the SNR to be enhanced in the presence of noise correlations in the NERM model, at least for large enough ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$. In essence, for the SNR to increase in the presence of noise correlations, it is necessary that the numerator $\Delta{r_{\rm R}}$ increases (which is not sufficient, since the denominator could increase too much). Furthermore, a necessary condition for $\Delta{r_{\rm R}}$ to increase is, in its turn, that ${r_{\rm R}^{(+)}}$ increases. Therefore, for the SNR to increase with ${\alpha_{\rm V}}>0$, ${r_{\rm R}^{(+)}}$ must increase, which implies ${\mu_{\rm I}}<{\mu_{\rm I}}^{*}$.
In Appendix J we also show that this upper bound to ${\mu_{\rm I}^{(+)}}<{\mu_{\rm I}}^{*}$ can be improved by taking into account the increase in the readout rate variance, i.e., by imposing that, beyond the numerator, the SNR itself should increase in the presence of noise correlations. With the analytic solution at hand, the condition $\text{SNR}_{\rm q}\geq\text{SNR}_{\rm w}$ implies that (see Appendix J) ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}<{\mu_{\rm I}}^{*}$, where $\hat{\mu_{\rm I}}$ is the solution of the transcendental equation $g_{1}^{\prime}+g_{1}^{2}=2{\tau_{\rm c}}{r_{\rm w}}g_{1}^{2}$. Although $\hat{\mu_{\rm I}}$ has been derived for low enough values of ${v_{\rm c}}$ and ${\mu_{\rm I}^{(-)}}$ (see Appendix J), it actually works as an upper bound for the value of $\mu$ below which we observe the SNR enhancement in all our simulations –see the green vertical line in Fig. 10. See also Figs. 11,30,31 and Appendix O.
Again, this necessary condition is not sufficient to guarantee that $\text{SNR}_{\rm q}>\text{SNR}_{\rm w}$. Too high values of ${v_{\rm c}}$, or non-negligible values of ${\mu_{\rm I}}^{(-)}$, hence of ${r_{\rm R}}^{(-)}$, may make $\text{SNR}_{\rm q}<\text{SNR}_{\rm w}$. Let us, first, focus on the dependence on ${\mu_{\rm I}^{(-)}}$. Mind that, at least for low enough ${\sigma_{\rm c}}$, ${v_{\rm R}}^{(-)}$ is an increasing function of ${\mu_{\rm I}}^{(-)}$ (see Fig. 10) and, therefore, the SNR denominator enhancement increases with ${\mu_{\rm I}^{(-)}}$. The larger the value of ${\sigma_{\rm c}}$, the lower the values of the (leftmost) inflection points of ${r_{\rm R}}$ and ${v_{\rm R}}$ (see Fig. 10). This implies that the larger ${\sigma_{\rm c}}$, the lower the value that ${\mu_{\rm I}}^{(-)}$ should take for the SNR to increase (as can be seen in Fig. 30 in Appendix N).
@@
We now focus on the dependence on ${\sigma_{\rm c}}$. For large enough values of ${\sigma_{\rm c}}$ (or ${v_{\rm c}}$), the SNR denominator increases faster with ${\sigma_{\rm c}}$ than the numerator $\Delta{r_{\rm R}}$, because of the over-Poissonian correction to the variance, cf. the second term in Eq. (13a), or Eq. (110) in Appendix I. This is the origin of the non-monotonous dependence of the SNR on the noise amplitude (see as well Appendix F for an explicit analysis of this).
@@
In sum, in order for the SNR to increase in the presence of noise correlations and for large enough ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$, it is necessary that: (c1) ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$ is low enough, and (c2) both ${\sigma_{\rm c}}$ and ${\mu_{\rm I}}^{(-)}<\hat{\mu_{\rm I}}^{(-)}$ are low enough, with the upper bound $\hat{\mu_{\rm I}}^{(-)}$ being a decreasing function of ${\sigma_{\rm c}}$.
@@
Figure 10: Readout rate mean, variance and excess of SNR for negligible values of ${r_{\rm R}}^{(-)}$. (A) Mean firing rate ${r_{\rm R}}({\mu_{\rm I}})$ of the white-noise LIF (dashed line), and the quenched-noise colored-noise LIF ${r_{\rm q}}({\mu_{\rm I}})$ (continuous lines) for various values of ${\sigma_{\rm c}}$ (or ${v_{\rm c}}={\sigma_{\rm c}}^{2}/({\tau_{\rm c}})$, different curves). For each ${\mu_{\rm I}}$, the value of ${\sigma_{\rm I}}$ is set to ${\sigma_{\rm I}}=(w{\mu_{\rm I}})^{1/2}$, with $w=10^{-3}$. As in Fig. 6, the circle and error-bars indicate the inflection point and ${v_{\rm c}}^{1/2}$, while the black dotted line indicate the activation function of the deterministic LIF $r_{0}$. (B) A zoom of the panel A, in which we report as well the quenched-noise rate standard deviation, ${v_{\rm q}}^{1/2}$ (dashed-dotted lines). The vertical lines indicate the maximum of each ${r_{\rm q}}({\mu_{\rm I}})$ curve. (C) The difference between the SNR $[{r_{\rm q}}({\mu_{\rm I}})-{r_{\rm q}}({\mu_{\rm I}}^{(-)})]/[({v_{\rm q}}({\mu_{\rm I}})+{v_{\rm q}}({\mu_{\rm I}}^{(-)})/2]^{1/2}$ (taking a low ${\mu_{\rm I}}^{(-)}=260$, actually equivalent to neglect ${r_{\rm q}}^{(-)}$) in the presence of colored noise, minus the white-noise SNR, $=[{r_{\rm w}}({\mu_{\rm I}})-{r_{\rm w}}({\mu_{\rm I}}^{(-)})]/[({v_{\rm w}}({\mu_{\rm I}})+{v_{\rm w}}({\mu_{\rm I}}^{(-)}))/2]^{1/2}$, assuming a Poissonian rate variance, ${v_{\rm w}}={r_{\rm w}}$. The vertical dashed line corresponds to our upper bound $\hat{\mu_{\rm I}}$. Rates and variances are calculated with a ${\sigma_{\rm I}}^{2}=w{\mu_{\rm I}}$ that varies with ${\mu_{\rm I}}$ for fixed $w$. The rest of the model parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.0025$, $w=0.001$, ${\mu_{\rm I}}^{(-)}=260$.